# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v3

@@ Abstract
In this work, we demonstrate a general phenomenology realizing an RG as a projection. For the construction of this projection, a mathematical structure called ideal plays a fundamental role. An ideal is fundamentally noninvertible and nonabelian (or nongroup-like). Moreover, it is also different from a subring, which commonly appears in physics literature. In principle, one can construct (infinitely) many new transformation laws of anyons by considering the ideal decomposition of existing fusion rings. More precisely, we study anyon transformation law or ring homomorphism between anyons in conformal field theories connected by massless RG flows[86]. Surprisingly, even in the simplest flow, such as the $SU(2)_{1}\times SU(2)_{1}$ Wess-Zumino-Witten (WZW) model to the $SU(2)_{2}$ WZW model and the flow from the tricritical Ising CFT to the Ising CFT, negative and fractional fusion coefficient, and unusual (or nonintegrable) flows appear inevitably. The unusual fusion coefficient has appeared recently in [87, 88, 57, 89, 90], and the unusual flow seems to correspond to partially solvable systems in [91, 92, 93].
@@ II Substructure from homomorphism: Projection as renormalization group flow
In this section, we denote general symmetry-based aspects of massless and massive RG flows. We mainly discuss $1+1$ dimensional CFTs or corresponding $2+1$ dimensional TOs as examples, but one can apply the same strategy to general models in physics in higher space-time dimensions. Our argument is elementary in contemporary mathematics, so we list a classic textbook as a reference[98]. The general algebraic expression of the massless RG[86] in this work is the following ring homomorphism,
@@
where $\mathbf{A}$ is an algebra describing generalized symmetries (or integral of motions) or underlying fusion category of anyons, the lower index $(1)$ and $(2)$ label the ultra-violet (UV) and infra-red (IR) theories, respectively, and $a,b$ are the algebraic objects. The above conditions represent that the structure of fusion is compatible or preserved under the application of domain walls, and the theory is transformed smoothly (or masslessly) into another theory. The corresponding argument has appeared evidently in [32] based on the earlier work [99] and the folding trick [11]. When interpreting this homomorphism as a defect between two theories, it is referred to as the RG domain wall, and it is expected to describe a massless RG[86]. We also note that the categorical structure corresponding to the ring homomorphism, called a functor, has been studied as gapped (or symmetry-preserving) domain walls in [22, 23, 25, 27] with emphasis on the theoretical aspect of TQFTs or TOs. It should also be stressed that the underlying category theory should be generalized to $\mathbb{C}$-linear category, where $\mathbb{C}$ represents the complex number field, because the gapped domain wall can produce noninteger coefficients[87]. For readers interested in the underlying phenomenology of this phenomenon, we explain it in Appendix B and note the related work by the first author[100].
@@
We restrict our attention to abstract algebra, not to category theory, because the underlying category theories can be premodular fusion or superfusion categories, or even outside of them that are still under development. However, the basic fusion ring in this manuscript is well-defined and has been studied several times both in the mathematics and physics communities. In other words, our discussion can be a clue to constructing a new fundamental class of categories. The fusion rings with negative or fractional coefficients are the corresponding structure, and they appeared in the study of fermionic string theories [124] and more recently in [88, 57], for example. The corresponding fusion rings have been studied more generally by the first author in [89, 90] with a close connection to the simple current extension by Schellekens and the collaborators[72, 73, 74, 75, 76, 77, 78, 79].
@@
For the application of our method to the classification of the gapped phases, we remind that the massive RGs have a close connection to boundary conformal field theories (BCFTs)[165, 166, 167, 168, 169]. The correspondence has been studied in a wide variety of communities under various names as Moore-Seiberg data[170, 171, 172, 173], Li-Haldane conjecture[174, 167], and boson condensation[175, 21]. The combination of our algebraic arguments and symmetry analysis of boundary state in [176] will provide the general criteria to describe the general class of gapped phases. Related arguments can be seen in [177, 64, 65, 178, 179, 180, 181]. However, as the author has clarified in [68], the label of boundary state can be outside of the original symmetry $\mathbf{A}_{(1)}$ or $\mathbf{A}_{(2)}$. The corresponding boundary states are known as new boundary conditions in condensed matter[182, 183, 184] and symmetry-breaking boundary conditions in high energy or mathematical physics[185, 186, 187] because these boundary condition breaks extended bulk symmetry such as $W$ or Lie group symmetry in the bulk CFT (we note [188] as a useful reference). For the later convenience, we introduce the extended ring (or extended algebra) $\mathbf{A}^{\text{ex}}$ with $\mathbf{A}\subset\mathbf{A}^{\text{ex}}$ and assume that the underlying CFT Hamiltonians are symmetric under $\mathbf{A}^{\text{ex}}$ action but the Hilbert spaces are labeled by $\mathbf{A}^{\text{ex}}$. Then one can express the set $\mathbf{R}\subset\mathbf{A}^{\text{ex}}$ labelling $\mathbf{A^{\prime}}$ invariant module as follows,
@@
Hence, in this formalism, one can formulate $\mathbf{R}$ as a generalization of an ideal because $r\in\mathbf{R}$ can be outside of $\mathbf{A^{\prime}}$. We propose that the above subring invariant set governs the massive RG and their classifications. We remind that whereas the set is invariant under the action of the subring, the object will be noninvariant, and this is nothing but the SSB.
@@ II.1 Unbroken and emergent symmetry
where $Q_{a^{\prime}_{(1)}}$ is a symmetry operator (or integral of motions) corresponding to the element $a^{\prime}_{(1)}\in\mathbf{A}_{(1)}^{\text{ub}}$. Inversely, if the perturbation $H_{\text{per}}$ is given, one can determine possible RG flows by studying the maximal subalgebra structure which commutes with $H_{\text{per}}$ and its disjoint ideals. Related arguments based on group symmetry can be seen ubiquitously, and our argument can be considered as a generalization emphasizing the ideal structure. We note [191, 192] as references containing analysis of symmetry with ring structure and their action on bulk fields.
@@ II.2 Historical remarks on the projections in physics
Related to the argument in CFT in this manuscript, we also note the restriction of Hilbert space in quantum group symmetric models[197, 198, 199] which have a connection to the Dotsenko-Fateev Coulomb gas[200, 201] (or Feigin-Fuchs construction[202] based on Wakimoto free field representation[203]) and the massless RG flow[86]. More recently, the RG domain wall, which is in the main scope of this work, has been introduced in [31] and its general correspondence with existing coset CFTs has been proposed in [32]. Some of the properties of homomorphism, or the properties of symmetry-preserving domain wall, have already appeared in [32], and it is remarkable that this pioneering work appeared before the introduction of generalized symmetry[97]. Later, more general cases have been studied in [204, 205, 206, 207, 208] and more mathematical understanding has been developed in [131, 132, 120, 209, 67, 210, 211, 212]. The idea of RG domain wall has also been combined with the holography in [213, 214] and related bound of entropy of conformal interface has been discussed in [215]. More recently, the correspondence between symmetry-preserving domain wall[22, 23, 24, 25, 27] and RG domain wall has been realized in [210, 66, 87, 216, 68], but this is relatively less familiar in the fields.
@@ IV Construction of homomorphism
In this section, we study ring homomorphisms between fusion rings appearing in massless RG flow of conformal field theories[86]. We study the massless RG flow between ultraviolet (UV) CFT $\mathbf{B}$ to infrared (IR) CFT $\mathbf{B}^{\prime}$ where $\mathbf{B}$ and $\mathbf{B}^{\prime}$ represent nonchiral fusion ring or anyon algebra of spherical fusion category (SFC)[245, 246, 162]. Roughly speaking, the SFC corresponds to the bosonic modular invariant and conformal bootstrap technique[247]. We concentrate our attention on the algebraic relation of objects or anyons, study the ring homomorphism $\rho$: $\mathbf{B}$ $\rightarrow$ $\mathbf{B}^{\prime}$ (or algebraic data of a functor between two SFCs[21, 23, 25]). To simplify the notation and discussion, we assume the theory is in the scope of Moore-Seiberg data[170, 171, 172, 173] and the ring homomorphism between nonchiral fusion rules naturally introduces a ring homomorphism between chiral fusion rules and vice versa. Algebraically, this is trivial because of the ring isomorphism between the nonchiral and chiral fusion rules.
@@ VI Acknowledgement
YF thanks Guangyue Ji, Ha-Quang Trung, and Bo Yang for the interesting discussions on the hierarchical structure of fractional quantum Hall systems and Taishi Kawamoto for the collaborations related to this project. YF also thanks Takamasa Ando, Weiguang Cao, and Pochung Chen for interesting discussions on noninvertible symmetries, Takumi Fukushima for sharing his knowledge of fracton topological orders, and Yifan Liu for reminding us of the significance of representing RG as a projection in a different research project, and Chihiro Matsui and Hosho Katsura for sharing their knowledge on the partial solvable systems and for the related discussions. YF thanks Sylvain Ribault for many useful comments on the earlier version of the draft. YF thanks Gerard Watts for pointing out the correspondence between the nontrivial homomorphism from the tricritical Ising CFT to the Ising CFT and phase diagram in [250] and Jurgen Fuchs for notifying us of terminology in category theory and for related useful comments. We also thank Yuya Kusuki for a useful comment. YF thanks the support from NCTS and CTC.
@@ Appendix B Conjecture on coset representation of symmetry preserving domain walls
Generalizing the argument in existing literature, we conjecture the decomposition of coset CFTs[111, 112, 113, 114]. The symbol “//" in this section is different from that of the quotient ring in the main text, so we placed the arguments here. As far as we have checked, the construction of ideals in the main text has not been achieved in general, and the study of the possible homomorphism between theories described by coset CFTs will be useful. The correspondence between coset conformal field theory and massless RG flow has been studied in [115], and its connection to current-current deformation has been studied in the context of $\lambda$ deformation[280, 281, 282, 283]. Related quantum wire junction problems have been studied in [284, 285, 286]. We also note more recent literature from different research motivations[287, 288]. In this section, we denote chiral or antichiral CFT as $M$ or $\overline{M}$ and label them by lower index IR or UV to identify them as UV and IR theories, respectively.
@@
where the product is a Deligne product and the object $M_{\text{UV}}\otimes_{\text{D}}M_{\text{IR}}$ appears in recent literature [289, 68].
By considering the corresponding quantum Hall point contact or quantum tri-wire junction problem[290], it is possible to interpret the decomposition more phenomenologically. The anomaly matching flow corresponds to the situation where the current flows from one wire to another wire with some dissipation at the junction. The anomaly cancellation flow corresponds to the case where the currents from one wire and another wire cancel at the junction point. In the main text, we have paid particular attention to the symmetry-preserving domain wall, but the coset representation will contain sectors corresponding to the charged domain wall. If there exists a correspondence between charged sectors and uncharged sectors both in UV and IR (or both systems have fractional supersymmetry), one can straightforwardly construct a mapping relation between UV and IR theories by shifting the charge appearing in the symmetry-preserving domain wall[68].
@@
$\mathbf{A}_{(2)}$ $\rightarrow\mathbf{A}_{(1)},$ (76)
$\rho(a_{(2)})$ $=\sum_{a_{\text{DW}},a_{(1)}}B^{\alpha_{(2)}}_{a_{\text{DW}},a_{(1)}}q_{a_{\text{DW}}}a_{(1)}$ (77)
@@
It should be stressed that even when assuming the NIM-rep of $B$, the resultant coefficient matrix $A$ can be outside of NIM-rep, because the quantum dimension can take noninteger values. Hence, the existence of the coset representation, or level-rank duality, inevitably induces the fusion ring homomorphisms with noninteger coefficients in general.
@@ Appendix C Noninteger coefficient in the sandwich construction
The right handsides perfectly reproduce the algebraic relation of the left hand sides, by identifying $\{I^{\prime},\psi^{\prime},e^{\prime},m^{\prime}\}$ as a subalgebra of $\{I,\psi,\overline{\psi},\epsilon,\sigma_{\text{Bulk},+},\mu_{\text{Bulk},+}\}$. This interpretation of the reduction by taking a subalgebra is also compatible with the intimate relationship between chiral CFT, BCFT, and bulk renormalization group in the literature. We also note related references studying subalgebra structure.

### v4

# Homomorphism, substructure, and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

@@ Abstract
In this work, we demonstrate a general phenomenology realizing an RG as a projection. For the construction of this projection, a mathematical structure called ideal plays a fundamental role. An ideal is fundamentally noninvertible and nonabelian (or nongroup-like). Moreover, it is also different from a subring, which commonly appears in physics literature. In principle, one can construct (infinitely) many new transformation laws of anyons by considering the ideal decomposition of existing fusion rings. More precisely, we study anyon transformation law or ring homomorphism between anyons in conformal field theories connected by massless RG flows[86, 87, 88]. Surprisingly, even in the simplest flow, such as the $SU(2)_{1}\times SU(2)_{1}$ Wess-Zumino-Witten (WZW) model to the $SU(2)_{2}$ WZW model and the flow from the tricritical Ising CFT to the Ising CFT, negative and fractional fusion coefficient, and unusual (or nonintegrable) flows appear inevitably. The unusual fusion coefficient has appeared recently in [89, 90, 57, 91, 92], and the unusual flow seems to correspond to partially solvable systems in [93, 94, 95].
@@ II Substructure from homomorphism: Projection as renormalization group flow
In this section, we denote general symmetry-based aspects of massless and massive RG flows. We mainly discuss $1+1$ dimensional CFTs or corresponding $2+1$ dimensional TOs as examples, but one can apply the same strategy to general models in physics in higher space-time dimensions. Our argument is elementary in contemporary mathematics, so we list a classic textbook as a reference[100]. The general algebraic expression of the massless RG[86, 87, 88] in this work is the following ring homomorphism,
@@
where $\mathbf{A}$ is an algebra describing generalized symmetries (or integral of motions) or underlying fusion category of anyons, the lower index $(1)$ and $(2)$ label the ultra-violet (UV) and infra-red (IR) theories, respectively, and $a,b$ are the algebraic objects. The above conditions represent that the structure of fusion is compatible or preserved under the application of domain walls, and the theory is transformed smoothly (or masslessly) into another theory. The corresponding argument has appeared evidently in [32] based on the earlier work [101] and the folding trick [11]. When interpreting this homomorphism as a defect between two theories, it is referred to as the RG domain wall, and it is expected to describe a massless RG[86, 87, 88]. We also note that the categorical structure corresponding to the ring homomorphism, called a functor, has been studied as gapped (or symmetry-preserving) domain walls in [22, 23, 25, 27] with emphasis on the theoretical aspect of TQFTs or TOs. It should also be stressed that the underlying category theory should be generalized to $\mathbb{C}$-linear category, where $\mathbb{C}$ represents the complex number field, because the gapped domain wall can produce noninteger coefficients[89]. For readers interested in the underlying phenomenology of this phenomenon, we explain it in Appendix B and note the related work by the first author[102].
@@
We restrict our attention to abstract algebra, not to category theory, because the underlying category theories can be premodular fusion or superfusion categories, or even outside of them that are still under development. However, the basic fusion ring in this manuscript is well-defined and has been studied several times both in the mathematics and physics communities. In other words, our discussion can be a clue to constructing an unexplored fundamental class of categories. The fusion rings with negative or fractional coefficients are the corresponding structure, and they appeared in the study of fermionic string theories [126] and more recently in [90, 57], for example. The corresponding fusion rings have been studied more generally by the first author in [91, 92] with a close connection to the simple current extension by Schellekens and the collaborators[72, 73, 74, 75, 76, 77, 78, 79].
@@
For the application of our method to the classification of the gapped phases, we remind that the massive RGs have a close connection to boundary conformal field theories (BCFTs)[167, 168, 169, 170, 171]. The correspondence has been studied in a wide variety of communities under various names as Moore-Seiberg data[172, 173, 174, 175], Li-Haldane conjecture[176, 169], and boson condensation[177, 21]. The combination of our algebraic arguments and symmetry analysis of boundary state in [178] will provide the general criteria to describe the general class of gapped phases. Related arguments can be seen in [179, 64, 65, 180, 181, 182, 183, 184]. However, as the author has clarified in [68], the label of boundary state can be outside of the original symmetry $\mathbf{A}_{(1)}$ or $\mathbf{A}_{(2)}$. The corresponding boundary states are known as new boundary conditions in condensed matter[185, 186, 187] and symmetry-breaking boundary conditions in high energy or mathematical physics[188, 189, 190] because these boundary condition breaks extended bulk symmetry such as $W$ or Lie group symmetry in the bulk CFT (we note [191] as a useful reference). For the later convenience, we introduce the extended ring (or extended algebra) $\mathbf{A}^{\text{ex}}$ with $\mathbf{A}\subset\mathbf{A}^{\text{ex}}$ and assume that the underlying CFT Hamiltonians are symmetric under $\mathbf{A}^{\text{ex}}$ action but the Hilbert spaces are labeled by $\mathbf{A}^{\text{ex}}$. Then one can express the set $\mathbf{R}\subset\mathbf{A}^{\text{ex}}$ labelling $\mathbf{A^{\prime}}$ invariant module as follows,
@@
Hence, in this formalism, one can formulate $\mathbf{R}$ as a generalization of an ideal because $r\in\mathbf{R}$ can be outside of $\mathbf{A^{\prime}}$. We propose that the above subring invariant set governs the massive RG and their classifications. We remind that whereas the set $\mathbf{R}$ is invariant under the action of the subring $\mathbf{A^{\prime}}$, the object $r\in\mathbf{R}$ will be noninvariant, and this is nothing but the SSB.
@@ II.1 Unbroken and emergent symmetry
where $Q_{a^{\prime}_{(1)}}$ is a symmetry operator (or integral of motions) corresponding to the element $a^{\prime}_{(1)}\in\mathbf{A}_{(1)}^{\text{ub}}$. Inversely, if the perturbation $H_{\text{per}}$ is given, one can determine possible RG flows by studying the maximal subalgebra structure which commutes with $H_{\text{per}}$ and its disjoint ideals. Related arguments based on group symmetry can be seen ubiquitously, and our argument can be considered as a generalization emphasizing the ideal structure. We note [194, 136, 137, 195, 196, 197, 198, 199] as references containing analysis of symmetry with ring structure and their action on bulk fields.
@@ II.2 Historical remarks on the projections in physics
Related to the argument in CFT in this manuscript, we also note the restriction of Hilbert space in quantum group symmetric models[204, 205, 206] which have a connection to the Dotsenko-Fateev Coulomb gas[207, 208] (or Feigin-Fuchs construction[209] based on Wakimoto free field representation[210]) and the massless RG flow[86, 87, 88]. More recently, the RG domain wall, which is in the main scope of this work, has been introduced in [31] and its general correspondence with existing coset CFTs has been proposed in [32]. Some of the properties of homomorphism, or the properties of symmetry-preserving domain wall, have already appeared in [32], and it is remarkable that this pioneering work appeared before the introduction of generalized symmetry[99]. Later, more general cases have been studied in [211, 212, 213, 214, 215] and more mathematical understanding has been developed in [133, 134, 122, 216, 67, 217, 218, 219]. The idea of RG domain wall has also been combined with the holography in [220, 221] and related bound of entropy of conformal interface has been discussed in [222]. More recently, the correspondence between symmetry-preserving domain wall[22, 23, 24, 25, 27] and RG domain wall has been realized in [217, 66, 89, 223, 68], but this is relatively less familiar in the fields.
@@ IV Construction of homomorphism
In this section, we study ring homomorphisms between fusion rings appearing in massless RG flow of conformal field theories[86, 87, 88]. We study the massless RG flow between ultraviolet (UV) CFT $\mathbf{B}$ to infrared (IR) CFT $\mathbf{B}^{\prime}$ where $\mathbf{B}$ and $\mathbf{B}^{\prime}$ represent nonchiral fusion ring or anyon algebra of spherical fusion category (SFC)[252, 253, 164]. Roughly speaking, the SFC corresponds to the bosonic modular invariant and conformal bootstrap technique[254]. We concentrate our attention on the algebraic relation of objects or anyons, study the ring homomorphism $\rho$: $\mathbf{B}$ $\rightarrow$ $\mathbf{B}^{\prime}$ (or algebraic data of a functor between two SFCs[21, 23, 25]). To simplify the notation and discussion, we assume the theory is in the scope of Moore-Seiberg data[172, 173, 174, 175] and the ring homomorphism between nonchiral fusion rules naturally introduces a ring homomorphism between chiral fusion rules and vice versa. Algebraically, this is trivial because of the ring isomorphism between the nonchiral and chiral fusion rules.
@@ VI Acknowledgement
The first author thanks Guangyue Ji, Ha-Quang Trung, and Bo Yang for the interesting discussions on the hierarchical structure of fractional quantum Hall systems and Taishi Kawamoto for the collaborations related to this project. The first author also thanks Takamasa Ando, Weiguang Cao, and Pochung Chen for interesting discussions on noninvertible symmetries, Takumi Fukushima for sharing his knowledge of fracton topological orders, and Yifan Liu for reminding us of the significance of representing RG as a projection in a different research project, and Chihiro Matsui and Hosho Katsura for sharing their knowledge on the partial solvable systems and for the related discussions. The first author thanks Sylvain Ribault for many useful comments on the earlier version of the draft. The first author thanks Gerard Watts for pointing out the correspondence between the nontrivial homomorphism from the tricritical Ising CFT to the Ising CFT and phase diagram in [257] and Jurgen Fuchs for notifying us of terminology in category theory and for related useful comments. We also thank Yuya Kusuki for a useful comment. The first author thanks the support from NCTS and CTC.
@@ Appendix B Conjecture on coset representation of symmetry preserving domain walls
Generalizing the argument in existing literature, we conjecture the decomposition of $UV$ theories by coset CFTs[113, 114, 115, 116] or that by level-rank dualities[287, 288, 289]. The symbol “//" in this section is different from that of the quotient ring in the main text, so we placed the arguments here. As far as we have checked, the construction of ideals in the main text has not been achieved in general, and the study of the possible homomorphism between theories described by coset CFTs will be useful. The correspondence between coset conformal field theory and massless RG flow has been studied in [117], and its connection to current-current deformation has been studied in the context of $\lambda$ deformation[290, 291, 292, 293]. Related quantum wire junction problems have been studied in [294, 295, 296]. We also note more recent literature from different research motivations[297, 298, 299]. In this section, we denote chiral or antichiral CFT as $M$ or $\overline{M}$ and label them by lower index IR or UV to identify them as UV and IR theories, respectively.
@@
where the product is a Deligne product and the object $M_{\text{UV}}\otimes_{\text{D}}M_{\text{IR}}$ appears in recent literature [196, 68]. The corresponding decomposition has been studied as the level-rank duality[287, 288, 289] and the method has applied to the $SU(N)$ fractional Hall states in [297].
By considering the corresponding quantum Hall point contact or quantum tri-wire junction problem[300], it is possible to interpret the decomposition more phenomenologically(Fig. 5). The anomaly matching flow corresponds to the situation where the current flows from one wire to another wire with some dissipation at the junction. The anomaly cancellation flow corresponds to the case where the currents from one wire and another wire cancel at the junction point. In the main text, we have paid particular attention to the symmetry-preserving domain wall, but the coset representation will contain sectors corresponding to the charged domain wall. If there exists a correspondence between charged sectors and uncharged sectors both in UV and IR (or both systems have fractional supersymmetry), one can straightforwardly construct a mapping relation between UV and IR theories by shifting the charge appearing in the symmetry-preserving domain wall[68].
@@
$\rho=I\otimes d:\mathbf{A}_{(1)}$ $\rightarrow\mathbf{A}_{(2)},$ (76)
$\rho(a_{(1)})$ $=\sum_{a_{\text{DW}},a_{(2)}}B^{\alpha_{(1)}}_{a_{\text{DW}},a_{(2)}}q_{a_{\text{DW}}}a_{(2)}.$ (77)
@@
It should be stressed that even when assuming the NIM-rep of $B$, the resultant coefficient matrix $A$ can be outside of NIM-rep, because the quantum dimension can take noninteger values. Hence, the existence of the coset representation, or level-rank duality[287, 288, 289], inevitably induces the fusion ring homomorphisms with noninteger coefficients in general. The hierarchical strucutre of related to the level-rank duality has been studied in [297], but this research direction is relatively less familiar.
@@ Appendix C Noninteger coefficient in the sandwich construction
The right handsides perfectly reproduce the algebraic relation of the left hand sides, by identifying $\{I^{\prime},\psi^{\prime},e^{\prime},m^{\prime}\}$ as a subalgebra of $\{I,\psi,\overline{\psi},\epsilon,\sigma_{\text{Bulk},+},\mu_{\text{Bulk},+}\}$. This interpretation of the reduction by taking a subalgebra is also compatible with the intimate relationship between chiral CFT, BCFT, and bulk renormalization group in the literature.